ClubChem

ClubChem is an undergraduate student organization that was founded around 1984 as an effort to establish a club that was open to any undergraduate who had an interest in chemistry. Funded solely by the MIT Chemistry department, ClubChem is currently administrated by an executive committee of about eight people and contains people from within and outside the course 5 department. There is no official ”membership”, per se, (and no fees either) because we want to welcome everyone who has an interest in chemistry to participate in any or all of our events. However, we do maintain a mailing list to keep people informed of upcoming events.

In recent years, ClubChem’s activities have been directed towards three main areas. First, our club encourages interaction between students and the course five faculty. The most frequent way we do this is to have professor dinners. About once a month (ideally), we invite a different professor to have dinner at a local restaurant. Anyone who wants to attend is welcome, and ClubChem picks up the check for the professor. (Everyone else has to pay for him/herself. Sorry. =P )

The second facet of our organization is involvement with the community. We go out to elementary schools to perform ”magic shows”, which consist of a few chemical tricks, including acid/base indicators, the iodine clock, nylon, and liquid nitrogen. Our intention is to introduce kids to the exciting world of chemistry, but we might begin to consider performing these shows for other audiences as well. If you are interested in performing some of these magic shows, please contact our vice-president, who is in charge of organizing the shows and the scheduling.

Finally, our club has a social purpose – to make the lives of ClubChem people a little better. We have study breaks (ideally once a month) which are open to everyone. The free food at the study breaks ranges from ice cream and cookies to Mexican to Chinese. (Of course, we’re always open to other suggestions.) In the past, we’ve held a barbeque and beach trips, and the idea of a ski trip has been mentioned on occasion. We also have t-shirts that we sell to undergrads and grad students alike. We currently do not have a publicity chair, so if announcing study breaks sounds like fun to you, let the exec members know at ccexec@mit.edu! =)

We tend to be a friendly bunch of people, so if you have any questions or comments about ClubChem, talk to one of the officers, or e-mail us at ccexec@mit.edu. Thanks for reading!

jklie@mit.edu, August 1999